OpenStudy (anonymous):

Help????????? \[\int\limits_{0}^{2}\sqrt{2x}(\sqrt{x}+\sqrt{5})dx\]

zepdrix (zepdrix):

Distributing sqrt(2x) gives us,\[\Large\rm \int\limits_0^2 \sqrt2 ~x+\sqrt{10}~x^{1/2}~~dx\]Then just integrate each term, applying the power rule for integration.

OpenStudy (anonymous):

first term becomes after integration \[\frac{ 1 }{ \sqrt{2} } x ^{2} + Constant\] do u know how to work out the second term?

OpenStudy (anonymous):

this might help you http://www.mathsisfun.com/calculus/integration-definite.html

OpenStudy (anonymous):

I'm not getting

OpenStudy (anonymous):

the given answers don't belong to this problem, may be they made a typo somewhere in the actual question??? ask for a second opinion please.

OpenStudy (anonymous):

@shubhamsrg help???

OpenStudy (shubhamsrg):

Yes, the answers don't match up with the questions.

OpenStudy (anonymous):

thanks @shubhamsrg :)

ganeshie8 (ganeshie8):

yeah the `c` in the options makes no sense as this is a definite integral

OpenStudy (anonymous):

thanks @ganeshie8

ganeshie8 (ganeshie8):

for that matter, even the variable also makes no sense..

ganeshie8 (ganeshie8):

it should evaluate to some number..

OpenStudy (anonymous):

yes yo are right I get it evaluating to 5.978

OpenStudy (anonymous):

@Fábio are you sure you've got the right question/answers above?

OpenStudy (anonymous):

Are online issues that I need to resolve this issue but could not get any results ... so I ask for help ..

OpenStudy (anonymous):

something is wrong with your online system then ;)

OpenStudy (anonymous):

got to go, see you later

OpenStudy (anonymous):

Ok, I will ask staff to review the issue. thank you